OpenStudy (anonymous):

The measure of an acute angle is less than 90 degrees. If an acute angle measures 3x - 15, then the inequality to represent its measure is 3x - 15 < 90. Solve to find the measure of x.

OpenStudy (anonymous):

is x<35

OpenStudy (anonymous):

The answer to the second question follows; 3x-15<90 3x+15=90 -15 -15 3x=75 /3 /3 x=25

OpenStudy (anonymous):

x>35